# Feature Intensity Mapping: Polycyclic Aromatic Hydrocarbon Emission from All Galaxies Across Cosmic Time

### v1

@@ Abstract
Line intensity mapping (LIM) is an emerging technique for probing the aggregate emission of a spectral line from all sources, without requiring individual detections. Through the wavelength–redshift relation, one can map the line-of-sight evolution of the line emission that traces the underlying large-scale structure in a spectral-imaging survey. In this work, we present a new technique—feature intensity mapping—as an extension of the LIM formalism to map broad spectral features in 3D, rather than the narrow emission lines typically targeted by LIM. By accounting for the convolution of spectral features with the instrument’s spectral response across redshift, our technique enables simultaneous constraints on the redshift-dependent emission from multiple features. This approach enables 3D intensity mapping with some of the brightest features in galaxies’ infrared spectra: the polycyclic aromatic hydrocarbon (PAH) emission bands. We forecast the detectability of PAH signals using feature intensity mapping with the ongoing SPHEREx mission in the near-infrared and the proposed PRIMA mission in the far-infrared. We find that of per redshift bin of widths and can be achieved at and with SPHEREx and PRIMA, respectively, for multiple PAH features, suggesting a promising prospect for mapping the aggregate PAH emission at cosmological distances with upcoming datasets.111ⓒ 2025. All rights reserved.
UTF8bsmi
@@ 1 Introduction
Line intensity mapping (LIM; for reviews, see E. D. Kovetz et al., 2017; J. L. Bernal & E. D. Kovetz, 2022) is a type of IM that targets specific spectral lines emitted by galaxies. Using narrow-band spectral imaging mapping, LIM probes the integrated line emission from all sources across the sky. Through the redshift–frequency relation, the three-dimensional (angular–spectral) LIM data cube directly traces the 3D spatial large-scale structure (LSS) of the Universe via cumulative line emission. LIM thus offers a powerful tool to study the 3D LSS, one of the most important cosmological observables, particularly at high redshift (), where traditional galaxy surveys face limitations. Moreover, different spectral lines are sensitive to different physical processes within the interstellar medium (ISM) and intergalactic medium (IGM). Performing LIM across multiple lines thus offers a unique opportunity to study galaxy formation and evolution in various phases across cosmic time (G. Sun et al., 2019).
@@
There are a number of ongoing and planned LIM experiments target a wide range of the electromagnetic spectrum: the most extensively studied regimes are the radio for the 21 cm line (for a review, see A. Liu & J. R. Shaw, 2020) and the (sub-)millimeter for C II and the CO rotational ladder (A. T. Crites et al., 2014; G. K. Keating et al., 2015, 2016; CONCERTO Collaboration et al., 2020; G. K. Keating et al., 2020; P. A. R. Ade et al., 2020; J. Vieira et al., 2020; K. S. Karkare et al., 2022; K. A. Cleary et al., 2022; CCAT-Prime Collaboration et al., 2023). In the infrared (IR) regime, the recently launched NASA SPHEREx mission (B. P. Crill et al., 2020) is carrying out an all-sky near-infrared (NIR) spectral survey, making it well-suited for LIM targeting multiple lines, including H, O III, H, and O II (Y.-T. Cheng et al., 2024), among others (R. M. Feder et al., 2024). For far-infrared (FIR) wavelengths, the proposed NASA Probe-class mission PRIMA (J. Glenn et al., in prep.) is well-suited for conducting such a large-area spectral-imaging survey. This would open new avenues for LIM using various FIR lines that trace star formation and black hole growth in galaxies (e.g., N II, H2, S III, and Si II; A. Moullet et al. (2023)).
@@
In addition to spectral lines, one of the most prominent features of the infrared spectral energy distribution (SED) of galaxies is the emission bands from polycyclic aromatic hydrocarbons (PAHs; A. Leger & J. L. Puget, 1984; L. J. Allamandola et al., 1985; J. D. T. Smith et al., 2004; M. W. Werner et al., 2004). This emission arises from a series of vibrational modes, spanning wavelengths from 3 m to 20 m, with typical spectral widths ranging from 0.1 to a few microns (J. D. T. Smith et al., 2007). PAH emission can contribute up to 20% of the total IR luminosity and serves as an important tracer of dust content in galaxies (J. D. T. Smith et al., 2007; T. S. Y. Lai et al., 2020), making PAH features well-suited targets for IM.
@@
The PAH emission spectrum of a galaxy encodes a wealth of information about its ISM and the processes that shape it, including star formation history. The overall abundance of PAHs is now known to evolve systematically with metallicity (e.g., C. W. Engelbracht et al., 2005; G. Aniano et al., 2020; T. S. Y. Lai et al., 2020; C. M. Whitcomb et al., 2024), hinting that gas-phase growth and destruction by ionizing photons may each play important roles in the PAH lifecycle (see discussion in C. M. Whitcomb et al., 2024). Indeed, PAHs are observed to be strongly depleted in H II regions (J. Chastenet et al., 2019; J. Sutter et al., 2024; J. Chastenet et al., 2025), with potential consequences for the PAH signal in compact, star-forming galaxies in the high-redshift Universe. The ratios of the feature strengths can indicate changes in size, charge state, and hydrogenation of PAHs, as well as the properties of the radiation field that illuminates them (e.g., A. G. G. M. Tielens, 2008; A. Maragkoudakis et al., 2020; B. T. Draine et al., 2021). The PAH emission features have been used as a tracer of star-formation (e.g., N. M. Förster Schreiber et al., 2004; E. Peeters et al., 2004; H. V. Shipley et al., 2016; T. S. Y. Lai et al., 2020), though it remains unclear with what fidelity PAH emission is correlated, or leads or lags, the evolution of the star-formation density with cosmic time. Certainly, correlations between PAH abundance and the specific star formation rate in galaxies have been observed (A. Rémy-Ruyer et al., 2015; A. Nersesian et al., 2019).
@@
Evidence has emerged that even the high-redshift Universe is enriched with PAHs. Emission in the 3.3 m PAH feature has been detected in a ultra-luminous infrared galaxy (ULIRG) with Spitzer, with suggestions of the feature in three additional ULIRGs (A. Sajina et al., 2009). A second high-redshift () detection of the 3.3 m feature was made in a lensed Lyman break galaxy with Spitzer (B. Siana et al., 2009), and recently JWST has observed the 3.3 m PAH feature in a lensed galaxy (J. S. Spilker et al., 2023). Moreover, the 2175 Å ultraviolet extinction feature, often attributed to PAHs (e.g., C. Joblin et al., 1992; W. W. Duley & S. Seahra, 1998; M. Steglich et al., 2010; Q. Lin et al., 2025), has now been observed in multiple galaxies (V. Markov et al., 2023; J. Witstok et al., 2023; R. Fisher et al., 2025). While JWST may be able to detect more bright PAH-emitting galaxies at cosmological redshifts, or detect high-redshift PAHs through rest-frame UV absorption in the 2175 Å feature, the vast majority of galaxies with PAHs will not be observed by JWST by virtue of being too faint, falling outside the small sky fraction JWST will ultimately observe at sufficient depth, and, at , radiating most of their PAH emission at wavelengths longer than JWST can observe.
@@
Applying IM techniques to the aggregate PAH emission background can complement observations on individual bright high-redshift galaxies. The aggregate background PAH emission has been detected through cross-correlations. For example, Y.-K. Chiang & B. Ménard (2019) detected cross correlation signals up to between the Wide-field Infrared Survey Explorer (WISE) 12 m map and spectroscopic galaxies from the Sloan Digital Sky Survey, and attributed the signal to the cosmic PAH background. R. Córdova Rosado et al. (2024) detected a small angular scale cross correlation signal between broad-band images from the WISE 12 m band and microwave temperature fluctuations measured by the Atacama Cosmology Telescope (ACT). They attributed the signal to the correlation between the cosmic infrared background in ACT and the PAH emission from unresolved galaxies in WISE.
@@
In this work, we introduce a new technique, which we refer to as “feature intensity mapping,” as an extension of LIM to broad spectral features such as PAHs in spectral-imaging surveys. By computing the convolution of the SED of a given feature with the observing filter profile as a function of redshift, our method models the full set of auto- and cross-power spectra between spectral channels. This framework captures both correlations from the same spectral feature across adjacent channels and those from different features tracing the same underlying large-scale structure. We then use these measurements to infer the redshift-dependent total emission of all features in a Bayesian framework. This technique adopts a similar spirit to the Bayesian Multi-line Intensity Mapping framework developed in Y.-T. Cheng et al. (2024).
The feature intensity mapping technique enables 3D intensity mapping of one of the brightest infrared SED features from PAHs, which would otherwise be inaccessible using the conventional LIM formalism. Furthermore, since PAHs are broad spectral features, the required spectral resolution is lower compared to LIM. This makes fast, wide-area mapping feasible even with low-resolution () spectral imaging. This technique is developed in time for the upcoming SPHEREx observations in the –m range, which will be sensitive to the 3.3 m PAH feature at (E. Zhang et al., 2025). The proposed PRIMA mission, targeting the – m wavelength range, is also well suited for mapping various PAH features across cosmic time (L. Bisigello et al., 2024, Faisst et al. in prep.). In this study, we apply our technique to forecast the sensitivity to the aggregate intensity of various PAH features for both SPHEREx and PRIMA. We find that high-significance detections, at the level of several tens of , for multiple PAH features can be achieved at for SPHEREx and for PRIMA. These forecasting results suggest the great promise of our feature intensity mapping technique for probing the 3D cosmic PAH background.
@@ 2.1 Intensity Field
where denotes the rest-frame wavelength222Throughout this manuscript, and refer to the observed wavelength and frequency, while the rest-frame wavelength and frequency are denoted by and , respectively., and the normalization factor is defined by , such that is normalized: .
@@
where is the comoving distance444The redshift and comoving distance are used interchangeably throughout this manuscript., is the luminosity function of feature at redshift and angular position , and and are the comoving angular diameter distance and luminosity distance, respectively. We define the following quantities to simplify the expression:
@@ 2.2 Power Spectrum
We measure the information from spectral-intensity maps in the spherical harmonics space, where the auto and cross angular power spectra characterize the covariance of the field. We focus on large-scale covariance on clustering scales, where the emission sources trace the underlying large-scale structure, and we neglect higher-order effects such as redshift-space distortions and lensing magnification. We also ignore fluctuations from nonlinear clustering and Poisson noise. Since large-scale fluctuations can be fully described by a Gaussian probability distribution, the ’s capture all two-point information on large scales. This framework follows previous work (Y.-T. Cheng et al., 2023, 2024).
@@ 3.1 SPHEREx
SPHEREx is conducting the first all-sky near-infrared spectro-imaging survey with a pixel size of , completing four full-sky passes over the nominal two-year mission duration. SPHEREx uses six broadband detector arrays spanning wavelengths from 0.75 to 5 m, with low-resolution spectroscopy enabled by linear variable filters (P. M. Korngut et al., 2018). The spectral coverage is divided into 17 channels per band, with varying spectral resolution: for bands 1–3 (covering 0.75–2.42 m, totaling 51 channels), for band 4 (2.42–3.82 m, 17 channels), for band 5 (3.82–4.42 m, 17 channels), and for band 6 (4.42–5.00 m, 17 channels).
Operating in a sun-synchronous orbit, SPHEREx visits the north and south ecliptic poles on every orbit, yielding integration time approximately 50 times deeper in those regions than the all-sky area (S. Bryan et al., in prep.). Consequently, SPHEREx will produce two deep-field mosaic maps of approximately 100 deg2 each, optimized for studying the extragalactic background light.
In this work, we consider spectral imaging maps from SPHEREx in both its all-sky and deep-field configurations. For the all-sky case, we adopt an observed sky fraction of to account for Galactic masking. For the deep fields, we consider a total survey area of 200 deg2. We use a Gaussian filter transmission profile with the FWHM given by the nominal spectral resolution. Figure 1 presents the spectral resolution and the rms noise levels for both the all-sky and deep-field modes across the 102 SPHEREx channels. The noise levels are derived from the SPHEREx Sky Simulation (B. P. Crill et al., 2025), which takes into account both instrumental noise and the photon noise from various sources of emission, with the zodiacal light being the dominant source of the photon noise.
@@ 3.2 PRIMA
The PRobe far-Infrared Mission for Astrophysics (PRIMA555https://prima.ipac.caltech.edu/, J. Glenn et al., in prep.) is a 1.8 m cryogenically cooled far-infrared space observatory, currently a proposed concept for the first NASA Probe-Class mission.
PRIMA is designed to include two instruments: an imaging instrument, PRIMAger (L. Ciesla et al., in prep.), and a spectrometer, FIRESS (C. M. Bradford et al., in prep.). PRIMAger consists of two hyperspectral imagers with low-resolution spectroscopy capability () enabled by linear variable filters, covering the wavelength range from 24 to 84 m, as well as four polarimetric imagers covering four broad bands between 92 and 235 m.
In this work, we focus on the spectral intensity maps observed by the two PRIMAger Hyperspectral Imagers, which we refer to as PHI hereafter. The PHI beam size ranges from to from short to long wavelengths. With a spectral resolving power of , there are 12 channels in total across the two hyperspectral imagers. We consider a fiducial PHI survey covering deg2 with a total integration time of 1000 hours. The instrument noise levels are derived from the official PRIMA website666https://prima.ipac.caltech.edu/page/instruments and shown in Figure 2.
@@
FIRESS provides two spectroscopy modes spanning 24 to 235 m. Its low-resolution mode uses four slit-fed grating spectrometer modules with spectral resolution , while a Fourier Transform Module can be inserted to provide high-resolution spectra with –. In this work, we focus on the low-resolution mode, assuming a uniform spectral resolution of , which yields a total of 229 spectral channels between 24 and 235 m. The FIRESS beam size varies from to across this wavelength range. We consider a fiducial FIRESS survey covering deg2 with an integration time of 1000 hours. The noise levels per spectral channel for FIRESS are also taken from the official PRIMA website and shown in Figure 2.
We adopt a multipole range of with 10 logarithmically spaced bins for both the PHI and FIRESS analyses. The higher range is selected to adapt to the small field sizes considered in these cases. We acknowledge that these higher modes probe scales where the signal enters the non-linear regime, beyond the validity of our current formalism. Modeling the effects of non-linear clustering is deferred to future work.

### v2

@@ Abstract
Line intensity mapping (LIM) is an emerging technique for probing the aggregate emission of a spectral line from all sources, without requiring individual detections. Through the wavelength–redshift relation, one can map the line-of-sight evolution of the line emission that traces the underlying large-scale structure in a spectral-imaging survey. In this work, we present a new technique—feature intensity mapping—as an extension of the LIM formalism to map broad spectral features in 3D, rather than the narrow emission lines typically targeted by LIM. By accounting for the convolution of spectral features with the instrument’s spectral response across redshift, our technique enables simultaneous constraints on the redshift-dependent emission from multiple features. This approach enables 3D intensity mapping with some of the brightest features in the infrared spectra of galaxies: the polycyclic aromatic hydrocarbon (PAH) emission bands. We forecast the detectability of PAH signals using feature intensity mapping with the ongoing SPHEREx mission in the near-infrared and the proposed PRIMA mission in the far-infrared. We find that of per redshift bin of widths and can be achieved at and with SPHEREx and PRIMA, respectively, for multiple PAH features, suggesting a promising prospect for mapping the aggregate PAH emission at cosmological distances with upcoming datasets.111ⓒ2025. All rights reserved.
@@ 1 Introduction
Line intensity mapping (LIM; for reviews, see E. D. Kovetz et al., 2017; J. L. Bernal & E. D. Kovetz, 2022) is a type of IM that targets specific spectral lines emitted by galaxies. Using narrow-band spectral-imaging mapping, LIM probes the integrated line emission from all sources across the sky. Through the redshift–frequency relation, the 3D (angular–spectral) LIM data cube directly traces the 3D spatial large-scale structure (LSS) of the Universe via cumulative line emission. LIM thus offers a powerful tool to study the 3D LSS, one of the most important cosmological observables, particularly at high redshift (), where traditional galaxy surveys face limitations. Moreover, different spectral lines are sensitive to different physical processes within the interstellar medium (ISM) and intergalactic medium (IGM). Performing LIM across multiple lines thus offers a unique opportunity to study galaxy formation and evolution in various phases across cosmic time (G. Sun et al., 2019).
@@
There are a number of ongoing and planned LIM experiments that target a wide range of the electromagnetic spectrum: the most extensively studied regimes are the radio for the 21 cm line (for a review, see A. Liu & J. R. Shaw, 2020) and the (sub)millimeter for C II and the CO rotational ladder (A. T. Crites et al., 2014; G. K. Keating et al., 2015, 2016; CONCERTO Collaboration et al., 2020; G. K. Keating et al., 2020; P. A. R. Ade et al., 2020; J. Vieira et al., 2020; K. S. Karkare et al., 2022; K. A. Cleary et al., 2022; CCAT-Prime Collaboration et al., 2023). In the infrared (IR) regime, the recently launched NASA SPHEREx mission (B. P. Crill et al., 2020) is carrying out an all-sky near-infrared (NIR) spectral survey, making it well suited for LIM targeting multiple lines, including H, O III, H, and O II (Y.-T. Cheng et al., 2024), among others (R. M. Feder et al., 2024). For far-infrared (FIR) wavelengths, the proposed NASA Probe-class mission concept—the Probe Far-infrared Mission for Astrophysics (PRIMA; J. Glenn et al., 2025)—is well suited for conducting such a large-area spectral-imaging survey. This would open new avenues for LIM using various FIR lines that trace star formation and black hole growth in galaxies (e.g., N II, H2, S III, and Si II; A. Moullet et al. (2023)).
@@
In addition to spectral lines, among the most prominent features of the IR spectral energy distribution (SED) of galaxies are the emission bands from polycyclic aromatic hydrocarbons (PAHs; A. Leger & J. L. Puget, 1984; L. J. Allamandola et al., 1985; J. D. T. Smith et al., 2004; M. W. Werner et al., 2004). Thes emissions arise from a series of vibrational modes, spanning wavelengths from 3 m to 20 m, with typical spectral widths ranging from 0.1 to a few microns (J. D. T. Smith et al., 2007). PAH emission can contribute up to 20% of the total IR luminosity and serves as an important tracer of dust content in galaxies (J. D. T. Smith et al., 2007; T. S. Y. Lai et al., 2020), making PAH features well-suited targets for IM.
@@
The PAH emission spectrum of a galaxy encodes a wealth of information about its ISM and the processes that shape it, including star formation history. The overall abundance of PAHs is now known to evolve systematically with metallicity (e.g., C. W. Engelbracht et al., 2005; G. Aniano et al., 2020; T. S. Y. Lai et al., 2020; C. M. Whitcomb et al., 2024), hinting that gas-phase growth and destruction by ionizing photons may each play important roles in the PAH lifecycle (see discussion in C. M. Whitcomb et al., 2024). Indeed, PAHs are observed to be strongly depleted in H II regions (J. Chastenet et al., 2019; J. Sutter et al., 2024; J. Chastenet et al., 2025), with potential consequences for the PAH signal in compact, star-forming galaxies in the high-redshift Universe. The ratios of the feature strengths can indicate changes in size, charge state, and hydrogenation of PAHs, as well as the properties of the radiation field that illuminates them (e.g., A. G. G. M. Tielens, 2008; A. Maragkoudakis et al., 2020; B. T. Draine et al., 2021). The PAH emission features have been used as tracers of star-formation (e.g., N. M. Förster Schreiber et al., 2004; E. Peeters et al., 2004; H. V. Shipley et al., 2016; T. S. Y. Lai et al., 2020), though it remains unclear with what fidelity PAH emission is correlated, or leads or lags, the evolution of the star-formation density with cosmic time. Certainly, correlations between PAH abundance and the specific star formation rate in galaxies have been observed (A. Rémy-Ruyer et al., 2015; A. Nersesian et al., 2019).
@@
Evidence has emerged that even the high-redshift Universe is enriched with PAHs. Emission in the 3.3 m PAH feature has been detected in a ultraluminous infrared galaxy (ULIRG) with Spitzer, with suggestions of the feature in three additional ULIRGs (A. Sajina et al., 2009). A second high redshift () detection of the 3.3 m feature was made in a lensed Lyman break galaxy with Spitzer (B. Siana et al., 2009), and recently JWST has observed the 3.3 m PAH feature in a lensed galaxy (J. S. Spilker et al., 2023). Moreover, the 2175 Å ultraviolet extinction feature, often attributed to PAHs (e.g., C. Joblin et al., 1992; W. W. Duley & S. Seahra, 1998; M. Steglich et al., 2010; Q. Lin et al., 2025), has now been observed in multiple galaxies (V. Markov et al., 2023; J. Witstok et al., 2023; R. Fisher et al., 2025). While JWST may be able to detect more bright PAH-emitting galaxies at cosmological redshifts, or detect high-redshift PAHs through rest-frame ultraviolet absorption in the 2175 Å feature, the vast majority of galaxies with PAHs will not be observed by JWST by virtue of being too faint, falling outside the small sky fraction JWST will ultimately observe at sufficient depth, and, at , radiating most of their PAH emission at wavelengths longer than JWST can observe.
@@
Applying IM techniques to the aggregate PAH emission background can complement observations on individual bright high-redshift galaxies. The aggregate background PAH emission has been detected through cross correlations. For example, Y.-K. Chiang & B. Ménard (2019) detected cross-correlation signals up to between the Wide-field Infrared Survey Explorer (WISE) 12 m map and spectroscopic galaxies from the Sloan Digital Sky Survey, and attributed the signal to the cosmic PAH background. R. Córdova Rosado et al. (2024) detected a small angular scale cross correlation signal between broadband images from the WISE 12 m band and microwave temperature fluctuations measured by the Atacama Cosmology Telescope (ACT). They attributed the signal to the correlation between the cosmic infrared background in ACT and the PAH emission from unresolved galaxies in WISE.
@@
In this work, we introduce a new technique, which we refer to as “feature intensity mapping,” as an extension of LIM to broad spectral features such as PAHs in spectral-imaging surveys. By computing the convolution of the SED of a given feature with the observing filter profile as a function of redshift, our method models the full set of auto- and cross-power spectra between spectral channels. This framework captures both correlations from the same spectral feature across adjacent channels and those from different features tracing the same underlying LSS. We then use these measurements to infer the redshift-dependent total emission of all features in a Bayesian framework. This technique adopts a similar spirit to the Bayesian Multi-line Intensity Mapping framework developed in Y.-T. Cheng et al. (2024).
The feature intensity mapping technique enables 3D IM of one of the brightest IR SED features from PAHs, which would otherwise be inaccessible using the conventional LIM formalism. Furthermore, since PAHs are broad spectral features, the required spectral resolution is lower compared to LIM. This makes fast, wide-area mapping feasible even with low-resolution () spectral imaging. This technique is developed in time for the upcoming SPHEREx observations in the –m range, which will be sensitive to the 3.3 m PAH feature at (E. Zhang et al., 2025). The proposed PRIMA mission, targeting the – m wavelength range, is also well suited for mapping various PAH features across cosmic time (L. Bisigello et al., 2024; A. L. Faisst et al., 2025). In this study, we apply our technique to forecast the sensitivity to the aggregate intensity of various PAH features for both SPHEREx and PRIMA. We find that high-significance detections, at the level of several tens of , for multiple PAH features can be achieved at for SPHEREx and for PRIMA. These forecasting results suggest the great promise of our feature intensity mapping technique for probing the 3D cosmic PAH background.
@@ 2.1 Intensity Field
where denotes the rest-frame wavelength,222Throughout this manuscript, and refer to the observed wavelength and frequency, while the rest-frame wavelength and frequency are denoted by and , respectively. and the normalization factor is defined by , such that is normalized: .
@@
where is the comoving distance,444The redshift and comoving distance are used interchangeably throughout this manuscript. is the luminosity function of feature at redshift and angular position , and and are the comoving angular diameter distance and luminosity distance, respectively. We define the following quantities to simplify the expression:
@@ 2.2 Power Spectrum
We measure the information from spectral-intensity maps in the spherical harmonics space, where the auto- and cross-angular power spectra characterize the covariance of the field. We focus on large-scale covariance on clustering scales, where the emission sources trace the underlying LSS, and we neglect higher-order effects such as redshift-space distortions and lensing magnification. We also ignore fluctuations from nonlinear clustering and Poisson noise. Since large-scale fluctuations can be fully described by a Gaussian probability distribution, the ’s capture all two-point information on large scales. This framework follows previous work (Y.-T. Cheng et al., 2023, 2024).
@@ 3.1 SPHEREx
SPHEREx is conducting the first all-sky NIR spectro-imaging survey with a pixel size of , completing four full-sky passes over the nominal two-year mission duration. SPHEREx uses six broadband detector arrays spanning wavelengths from 0.75 to 5 m, with low-resolution spectroscopy enabled by linear variable filters (P. M. Korngut et al., 2018). The spectral coverage is divided into 17 channels per band, with varying spectral resolution: for bands 1–3 (covering 0.75–2.42 m, totaling 51 channels), for band 4 (2.42–3.82 m, 17 channels), for band 5 (3.82–4.42 m, 17 channels), and for band 6 (4.42–5.00 m, 17 channels).
Operating in a sun-synchronous orbit, SPHEREx visits the north and south ecliptic poles on every orbit, yielding integration time approximately 50 times deeper in those regions than the all-sky area (S. Bryan et al., 2025). Consequently, SPHEREx will produce two deep-field mosaic maps of approximately 100 deg2 each, optimized for studying the extragalactic background light.
In this work, we consider spectral-imaging maps from SPHEREx in both its all-sky and deep-field configurations. For the all-sky case, we adopt an observed sky fraction of to account for Galactic masking. For the deep fields, we consider a total survey area of 200 deg2. We use a Gaussian filter transmission profile with the FWHM given by the nominal spectral resolution. Figure 1 presents the spectral resolution and the rms noise levels for both the all-sky and deep-field modes across the 102 SPHEREx channels. The noise levels are derived from the SPHEREx Sky Simulation (B. P. Crill et al., 2025), which takes into account both instrumental noise and the photon noise from various sources of emission, with the zodiacal light being the dominant source of the photon noise.
@@ 3.2 PRIMA
The PRobe far-Infrared Mission for Astrophysics (PRIMA;555https://prima.ipac.caltech.edu/ J. Glenn et al., 2025) is a 1.8 m cryogenically cooled far-infrared space observatory, currently a proposed concept for the first NASA Probe-Class mission.
PRIMA is designed to include two instruments: an imaging instrument, PRIMAger (L. Ciesla et al., 2025), and a spectrometer, FIRESS (C. M. M. Bradford et al., 2025). PRIMAger consists of two hyperspectral imagers with low-resolution spectroscopy capability () enabled by linear variable filters, covering the wavelength range from 24 to 84 m, as well as four polarimetric imagers covering four broad bands between 92 and 235 m.
In this work, we focus on the spectral-intensity maps observed by the two PRIMAger Hyperspectral Imagers (which we refer to as PHI hereafter). The PHI beam size ranges from to from short to long wavelengths. With a spectral resolving power of , there are 12 channels in total across the two hyperspectral imagers. We consider a fiducial PHI survey covering deg2 with a total integration time of 1000 hr. The instrument noise levels are derived from the official PRIMA website666https://prima.ipac.caltech.edu/page/instruments and shown in Figure 2.
@@
FIRESS provides two spectroscopy modes spanning 24 to 235 m. Its low-resolution mode uses four slit-fed grating spectrometer modules with spectral resolution , while a Fourier Transform Module can be inserted to provide high-resolution spectra with –. In this work, we focus on the low-resolution mode, assuming a uniform spectral resolution of , which yields a total of 229 spectral channels between 24 and 235 m. The FIRESS beam size varies from to across this wavelength range. We consider a fiducial FIRESS survey covering deg2 with an integration time of 1000 hr. The noise levels per spectral channel for FIRESS are also taken from the official PRIMA website and shown in Figure 2.
We adopt a multipole range of with 10 logarithmically spaced bins for both the PHI and FIRESS analyses. The higher range is selected to adapt to the small field sizes considered in these cases. We acknowledge that these higher modes probe scales where the signal enters the nonlinear regime, beyond the validity of our current formalism. Modeling the effects of nonlinear clustering is deferred to future work.